A wet porus substance in the open air loses its moisture at a rate proportional to the
moisture content, if a sheet hung in the wind loses half its moisture during the first
hour, then find the time when it has lost 95% moisture provided the weather
conditions remain the same.
1
Expert's answer
2020-03-17T13:56:59-0400

Let the moisture content be m.Thus, rate of moisture loss is given by "-dm\/dt" .

Given; "-{dm/dt} \propto m \implies dm/dt = -\lambda m" , where "\lambda" is a constant.

"\int dm/m =-\lambda \int dt"

"\implies m=m_0e^{- \lambda t}" ; where "m_0" is the initial moisture content at "t=0" .

Also given, "m_0/2=m_0e^{-\lambda} \implies \lambda =(ln2 )hr^{-1}"

If 95% moisture is lost at time t; then "m=5\%(m_0)=0.05m_0"

Thus, "0.05 m_0=m_0e^{-t*ln2} \implies t=log_2(20)=4.322hr.=259.3min."

Hence, the time when it has lost 95% moisture provided the weather conditions remain the same is after 4.322 hours.